OpenStudy (anonymous):

multiply the polynomials using distributive property and combine like terms: (6x+5)(2x-1)

OpenStudy (amistre64):

or... simply multipy them using multiplication :)

OpenStudy (anonymous):

6x*2x=12x^2 6x*-1=-6x 5*2x=10x 5*-1=-5 12x^2-6x+10x-5\[12x^2+4x-5\]

OpenStudy (amistre64):

6x +5 x 2x -1 --------- - 6x -5 12x^2 +10x ----------------- ?? ?? ??

OpenStudy (anonymous):

Who said multiplication wasn't addition....

OpenStudy (amistre64):

some german guy exiled to france for sayingit :)